Bravissimi, Inc. is going to pay a dividend of $4, and the dividends are expected to grow at a constant rate of 5% every year. The beta of Bravissimi's stock will be constant at 0.62. The T-bill rate is currently estimated to be 5%, and the return of S&P500 index is expected to be 12%.

a. What rate of return Bravissimi's investors require? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Market Capitalization rate = ?



b. Calculate the current value of Bravissimi's stock. (Do not round intermediate calculations. Round your answer to 2 decimal places.)

Intrinsic Value = ?
Answer #1

Market capitalisation rate = ke = 0.05+(0.62*(0.12-0.05)) D1 = 9.34% 5.00% $4.000 $92.17 4/(9.34% - 5%) Price = 01/(ke-g)